OpenStudy (anonymous):

Solve 2 cot^2 x - cot x = 3 at interval (0, 2pi)

zepdrix (zepdrix):

Move the 3 to the other side, then replace COT X with something simple.. like U. And then look at it, and think back to algebra, how you would solve a problem like that.

OpenStudy (anonymous):

Let y = cot x. Substitution! Solve for y from the quadratic formula and then put that back into y = cot x.

OpenStudy (anonymous):

@zepdrix , multiply the other side by 3?

zepdrix (zepdrix):

subtract 3 from both sides :D

OpenStudy (anonymous):

If you use what zepdrix and I are saying, you will get: \[ 2y^2 - y - 3 = 0. \]Once you solve for that. You can find the arccot of that, to get your answer.

OpenStudy (anonymous):

@Ahaanomegas, using the quadratic formula?

OpenStudy (anonymous):

Yes, use quadrant formula.

OpenStudy (anonymous):

I got 3/2 and -1

OpenStudy (anonymous):

Well, I got x = \[x = \tan^{-1} \frac{4}{1 \pm 5} = 33.69º \text{ and }\frac{3 \pi}{4}\]

OpenStudy (anonymous):

Correction: Quadratic formula. Not quadrant. ;)